OpenStudy (anonymous):

A large emerald with a mass of 378.24 grams was recently discovered in a mine. If the density of the emerald is 3.91grams over centimeters squared, what is the volume? Round to the nearest hundredth when necessary and only enter numerical values, which can include a decimal point.

OpenStudy (anonymous):

@phi can you help?

OpenStudy (anonymous):

@jim_thompson5910 ?

jimthompson5910 (jim_thompson5910):

The density formula is D = M/V where D = density M = mass V = volume

jimthompson5910 (jim_thompson5910):

what two things do we know?

OpenStudy (anonymous):

well i know the mass is 378.24

jimthompson5910 (jim_thompson5910):

what else does it give you

OpenStudy (anonymous):

3.91\[\frac{ grams }{ cm ^{2} }\]

jimthompson5910 (jim_thompson5910):

so D = 3.91 M = 378.24 V = unknown (leave it as V for now)

jimthompson5910 (jim_thompson5910):

plug them into the given formula to get D = M/V 3.91 = 378.24/V now solve for V

OpenStudy (anonymous):

okay, so i would multiply by 378.24?

jimthompson5910 (jim_thompson5910):

first you multiply both sides by V to get 3.91 = 378.24/V 3.91V = 378.24

jimthompson5910 (jim_thompson5910):

then you divide both sides by 3.91 to completely isolate V

OpenStudy (anonymous):

oh okay.

OpenStudy (anonymous):

96.73

jimthompson5910 (jim_thompson5910):

what are you getting after the 3?

jimthompson5910 (jim_thompson5910):

give me 6 decimal places of what you're getting

OpenStudy (anonymous):

oh my bad. 96.74

jimthompson5910 (jim_thompson5910):

good, you rounded properly this time You'll get V = 96.7365728900256 which rounds to V = 96.74

OpenStudy (anonymous):

i got it....thank you!

jimthompson5910 (jim_thompson5910):

yw